What is the vapor pressure in mmHg of a solution made by dissolving 24.1 g of NaCl in 500.0 g of water at 70°C, assuming a van't Hoff factor of 1.9? The vapor pressure of pure water at 70°C is 233.7 mmHg.
1
Expert's answer
2018-02-02T05:02:08-0500
Raoult's law: p = pH2O•x
n(NaCl)= 24.1/58.5 = 0.41 mol
n(H2O) = 500/18 = 27.78 mol
x(H2O) = 27.78/28.19 = 0.9855
p = 233.7•0.9855 =230.31 mmHg
p(exp) = i•p(teor) = 1.9•230.31 = 437.59 mmHg